OpenStudy (anonymous):

decide wheteher the pair of lines is parellel, perpendicular, or neither 2x+3y=1 3x =2y=7

OpenStudy (aroub):

3y=-2x+1 y=-2/3x+1/3 2y=-3x+7 y=-3/2+7/2 so i guess neither / wait is that right 3x=2y=7 or its a typing mistake?

OpenStudy (anonymous):

yeah that is right it is not a typing mistake

OpenStudy (aroub):

oh if its right then the answer is the lines are pperependicular

OpenStudy (anonymous):

The answer is neither. I have posted your question and explanation at: http://www.oojih.com/show/algebra/linearfunction/ scroll down to comment section to see your question. cheers

OpenStudy (anonymous):

Oops, I stand corrected, the answer is perpendicular.